OpenStudy (anonymous):

Medal Write as a fraction. 31 1/4%

OpenStudy (wolf1728):

per cent means for every hundred So for every hundred, you have 31.25 "things" So the ratio of "things" to 100 is 31.25 / 100

OpenStudy (anonymous):

so what do you do next

OpenStudy (wolf1728):

That is the fraction and the answer

OpenStudy (anonymous):

but how can you have a decimal

OpenStudy (wolf1728):

okay if you want an actual fraction .3125 / 1 OR 3,125 / 10,000

OpenStudy (anonymous):

like 6 1/4 % as a fraction is 1/16

OpenStudy (texaschic101):

3125/10000 reduces to 5/16

OpenStudy (wolf1728):

okay - I forgot to reduce the fraction

OpenStudy (texaschic101):

6.25% = 625/10000 = 1/16

OpenStudy (anonymous):

so the answer is 5/16?

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

than you everyone :)

OpenStudy (texaschic101):

:)

OpenStudy (wolf1728):

u r welcome

OpenStudy (anonymous):

i am not sure who to give the medal lol

OpenStudy (anonymous):

because you both helped me

OpenStudy (texaschic101):

give to wolf

OpenStudy (anonymous):

ok :)

OpenStudy (wolf1728):

I gave a medal to texaschic

OpenStudy (texaschic101):

thx...we're all good now :) everyone's happy

OpenStudy (wolf1728):

okay to us all